OpenStudy (anonymous):

What is the linear funtion of the graph with the points (1960,215) and (1955,210)? if you need more infor just ask its a line of best fit but for some reason i keep forgetting how to do it.

OpenStudy (anonymous):

slope is \[\frac{215-210}{1960-1955}=\frac{5}{5}=1\]

OpenStudy (anonymous):

okay i kept forgetting which was over which but i cant remeber anything for the b value and how to caculate it

OpenStudy (anonymous):

then use point slope formula to get \[y-215=x-1960\] \[y=x-1745\]

OpenStudy (anonymous):

formula is \[y-y_1=m(x-x_1)\] but easy in this case because m = 1

OpenStudy (anonymous):

okay thanks i new i was forgetting some formula